# The Narrow Emission Lines of Seyfert 1 Galaxies: Comparisons with a Large SDSS Sample

### v1

@@ 1 Introduction
The starting point for this paper is the dataset of SDSS spectra of AGN assembled and measured by one of us (H2005). Using the same procedures she had applied to the Data Release 2 of the SDSS in 2005, this new sample has included many more AGN, since it is based on all the emission line galaxy spectra available in Data Release 7. See §2 for an in-depth description of our dataset.
1.1 The nature of the narrow line emission from Seyfert 1 galaxies
@@
This paper is organized as follows. In the following section (§2) we describe our data and how it was processed. In section §3 we describe how the data sample was divided into several sub-groups, and investigate how these groups fall in the traditional line-ratio diagrams as well as in an alternate diagram introduced by L. J. Kewley et al. (2006). Section §4 discusses the prediction of [NII] from the [SII] doublet. In sections §5, 6, and 7 gas reddening, electron densities, and ionization parameters are estimated and compared, respectively. Sections §8 and 9 we analyze and interpret absorption line measurements of the starlight continuum in our categories of AGN spectra, and possible contamination from AGN components. Finally, in section §10 we discuss the implications of our results for selection of Seyfert galaxy samples, and the standard unification theory for Type I and Type II Seyferts. Our conclusions are summarized in §11.
2 The AGN and Star-Forming Galaxy Sample
@@
2.1 Hao-Sample Processing procedure
@@
2.2 Separation of the Narrow Line Region in the Broad Line Seyfert 1 Galaxies
@@
2.3 Seyfert 1 Sub-Classes from Relative Strengths of Broad and Narrow Balmer Lines
@@
2.4 Narrow-Line AGN Sample: Seyfert 2 and LINERs
@@
Unless a galaxy spectrum shows a 50% contribution from the LINER component, it cannot be reliably classified as LINER-dominated. Any emission-line galaxy lying above this equation 6 line is considered Seyfert 2-dominated. See §3.3 for further LINER details.
@@
Values of significantly above the right- hand side of equation 7 indicate that a classic AGN-powered NLR dominates the [OIII] emission. Our suspicion was that the Sy2U sub-class are more dominated by pure AGN NLR emission, whereas the Sy2L sub-class has more low-ionization ’contaminations’, which pull them down closer to the Ke01-line. Our subsequent analysis confirms and quantifies this supposition (see sections §5,§6, and §7 ). 666These two Sy2 sub-classes have the same average redshifts. Although, by design the Sy2U class of course has more powerful [OIII] emission than the Sy2L’s ( mean versus , both groups have the same mean H luminosities, .
2.5 Comparison Samples of Star-Forming Galaxies
@@
2.6 Stellar Masses
Although the detailed individual spectral modeling needed to calculate accurate stellar masses of our galaxies is beyond the scope of this work, we obtained rough estimates for each of the galaxy sub-groups in our study. We base these on the integrated i-band light given in the SDSS photometric database (https://www.sciserver.org/, using DR7 to match our spectroscopic data). Since the average redshifts of all of our AGN sub-groups are the same, it is not surprising that their average total i-band luminosities, assumed to be almost entirely starlight, are also very similar. In fact, the average absolute magnitude of all categories of Seyfert 1’s, Sy2’s and LINERs are within 0.1 mag of -21.6. Assuming the mass/light ratio of typical early type galaxies, of 2.5 in solar units, this corresponds to . We obtain similar stellar masses for our “Lower” and “Middle” SF galaxies, which have mean absolute i magnitude of -21.1 and -20.8 respectively. (The average absolute magnitude of our SF “Lower” galaxies, which we argue do not match the higher metallicities of Seyfert host galaxies, is -19.8).
3 Narrow Emission Line Ratios in Seyfert 1 Spectra
As noted in section §1.1, it has been widely assumed that any narrow lines present in a Seyfert 1 spectra are produced by the same ionization mechanism and originate from the same location as in a Sy2 AGN. In all BPT diagrams as well as in the Ke06 diagram, the Sy2s are clustered toward the upper right. Thus, if the NLR Sy1s are all AGN-powered, they should position themselves at the same location. But if these NLRs are contaminated by HII-regions or LINER gas, their positions in the diagrams will be shifted downwards.
@@ 3 Narrow Emission Line Ratios in Seyfert 1 Spectra
3.1 Shifts of Seyfert 1 sub-types in BPT Diagrams Caused by HII Regions
@@
3.2 Mixing Narrow Lines from Seyfert Nuclei and Star Formation: A simple model
@@ 3.2.1 Model Procedure
For the SF emission component we therefore take the mid-point values of the line-ratios between SF-Lower and SF-Middle as having 0% emissions from AGNs and 100% from HII-regions. For the pure AGN component, we select the median in the upper central part of the Sy2U region in all BPT diagrams (see section 2.4).
@@
Note. — Average Estimated percentage of AGN and HII Regions contribution in each galaxy group for our Toy Model. This simplified analysis excludes any possible LINER contribution (see section §3.3). The top column headers indicate the x-axis on the BPT diagrams, except for the right-most column header, which indicates the y-axis as adopted from L. J. Kewley et al. (2006). Ke01 and Kaf03 indicate the percentage of the type of galaxies laying above or below these boundary lines in the BPT diagrams.
3.3 LINER Components
@@
4 Predicting [NII] Flux from [SII] to Deblend Halpha
@@
5 Reddening Estimates from Narrow Emission Line Ratios
@@ 5.1 Milky Way Reddening
5.2 Average Reddenings of AGN Categories
@@
5.3 Comparison of Balmer Decrements with [SII]/[OII] Ratios
@@
6 Electron Densities in Different Galaxy Classes
@@
7 Are Weak Seyfert 1’s Explained by Low Ionization Parameter, Rather than Star Formation Contamination?
@@
8 Absorption Line Measurements of Stellar Populations in Seyfert Galaxies
@@
9 Small Contamination from Seyfert 1 Nonstellar Light
@@
The dashed line in each of the Figures 11 and 12 is a simple straight-line fit for pure starlight being emitted from pure stellar populations in a galaxy, from the top where SF-Lower galaxies are located, down to the LINERs in the lower right of these figures. This sequence is based on the fact that the continuum in star-forming galaxies is dominated by light from hot, luminous young stars, whereas LINER galaxies contain an older stellar population with no star formation in the last Gyr.
@@
We searched further for evidence of nonstellar AGN light by examining two additional measures of the continuum which are sensitive to violet light, where the AGN contribution is strongest. We show in Figure 12b the integrated color of the entire galaxy, obtained directly from the SDSS SkyServer photometry catalog, averaged over each galaxy class. In Figure 11, both and D4000 are compared against our stellar population age indicator, H absorption.
The same stellar population age sequence is evident in Figure 12 as we found in Figure 11. Again, the color and especially D4000 become smaller in the less AGN-dominated spectra. In other words, the intermediate Sy1’s, the Sy2L’s, and especially the weak Sy1’s show this indication of a larger proportion of blue continuum light from younger stars. Their average stellar populations are significantly younger than those of Passive Galaxies, LINERs, and Sy2U’s.
However, there is an additional effect evident in the color index: the strong Sy1’s (Sy1.2 and especially Sy1.0) shift dramatically to the blue (left), even though their weak H absorption indicates few young stars.
Thus the strong Seyfert 1’s have substantial extra flux in the u bandpass, which is not produced by hot stars. Instead, we attribute this extra u light to strong BLR emission contamination. Indeed, M. A. Malkan & W. L. W. Sargent (1982) first measured the BLR contributions produced by the blended broad-lines of the higher Balmer series above H, which then merge into a pseudo-continuum rising smoothly up to 3646Å, the onset of the Balmer Continuum emission jump. They found that the Balmer jump can produce as much flux at 3646Å as the underlying continuum. Based on their finding that the Balmer continuum flux is relatively flat throughout most of the sensitive part of the u filter bandpass (3350–3750Å), the typical Sy1 studied by M. A. Malkan & W. L. W. Sargent (1982) would have its integrated u flux enhanced by a factor of roughly 50%. The horizontal vector shows the effect of this BLR contamination on shifting the color of strong Sy1’s up to 0.44 mag to the left. In the intermediate and weak Seyfert 1’s, with substantially less broad H emission, this BLR contamination is too small to be measurable.
@@
10 Selection Effects in Seyfert Galaxy Surveys
@@
10.1 “Strong Unification” of Seyfert 1 and 2 Galaxies
@@
10.2 Are All Broad Line Regions the Same?

### v2

@@ I Introduction
The starting point for this paper is the dataset of SDSS spectra of AGN assembled and measured by one of us (H2005). Using the same procedures she had applied to the Data Release 2 of the SDSS in 2005, this new sample has included many more AGN, since it is based on all the emission line galaxy spectra available in Data Release 7. See §II for an in-depth description of our dataset.
I.1 The nature of the narrow line emission from Seyfert 1 galaxies
@@
This paper is organized as follows. In the following section (§II) we describe our data and how it was processed. In section §III we describe how the data sample was divided into several sub-groups, and investigate how these groups fall in the traditional line-ratio diagrams as well as in an alternate diagram introduced by L. J. Kewley et al. (2006). Section §IV discusses the prediction of [NII] from the [SII] doublet. In sections §V, VI, and VII gas reddening, electron densities, and ionization parameters are estimated and compared, respectively. Sections §VIII and IX we analyze and interpret absorption line measurements of the starlight continuum in our categories of AGN spectra, and possible contamination from AGN components. Finally, in section §X we discuss the implications of our results for selection of Seyfert galaxy samples, and the standard unification theory for Type I and Type II Seyferts. Our conclusions are summarized in §XI.
II The AGN and Star-Forming Galaxy Sample
@@
II.1 Hao-Sample Processing procedure
@@
II.2 Separation of the Narrow Line Region in the Broad Line Seyfert 1 Galaxies
@@
II.3 Seyfert 1 Sub-Classes from Relative Strengths of Broad and Narrow Balmer Lines
@@
II.4 Narrow-Line AGN Sample: Seyfert 2 and LINERs
@@
Unless a galaxy spectrum shows a 50% contribution from the LINER component, it cannot be reliably classified as LINER-dominated. Any emission-line galaxy lying above this equation 6 line is considered Seyfert 2-dominated. See §III.3 for further LINER details.
@@
Values of significantly above the right- hand side of equation 7 indicate that a classic AGN-powered NLR dominates the [OIII] emission. Our suspicion was that the Sy2U sub-class are more dominated by pure AGN NLR emission, whereas the Sy2L sub-class has more low-ionization ’contaminations’, which pull them down closer to the Ke01-line. Our subsequent analysis confirms and quantifies this supposition (see sections §V,§VI, and §VII ). 666These two Sy2 sub-classes have the same average redshifts. Although, by design the Sy2U class of course has more powerful [OIII] emission than the Sy2L’s ( mean versus , both groups have the same mean H luminosities, .
II.5 Comparison Samples of Star-Forming Galaxies
@@
III Narrow Emission Line Ratios in Seyfert 1 Spectra
As noted in section §I.1, it has been widely assumed that any narrow lines present in a Seyfert 1 spectra are produced by the same ionization mechanism and originate from the same location as in a Sy2 AGN. In all BPT diagrams as well as in the Ke06 diagram, the Sy2s are clustered toward the upper right. Thus, if the NLR Sy1s are all AGN-powered, they should position themselves at the same location. But if these NLRs are contaminated by HII-regions or LINER gas, their positions in the diagrams will be shifted downwards.
@@ III Narrow Emission Line Ratios in Seyfert 1 Spectra
III.1 Shifts of Seyfert 1 sub-types in BPT Diagrams Caused by HII Regions
@@
III.2 Mixing Narrow Lines from Seyfert Nuclei and Star Formation: A simple model
@@ III.2.1 Model Procedure
For the SF emission component we therefore take the mid-point values of the line-ratios between SF-Lower and SF-Middle as having 0% emissions from AGNs and 100% from HII-regions. For the pure AGN component, we select the median in the upper central part of the Sy2U region in all BPT diagrams (see section II.4).
@@
Note. — Average Estimated percentage of AGN and HII Regions contribution in each galaxy group for our Toy Model. This simplified analysis excludes any possible LINER contribution (see section §III.3). The top column headers indicate the x-axis on the BPT diagrams, except for the right-most column header, which indicates the y-axis as adopted from L. J. Kewley et al. (2006). Ke01 and Kaf03 indicate the percentage of the type of galaxies laying above or below these boundary lines in the BPT diagrams.
III.3 LINER Components
@@
III.4 Stellar Masses
Although the detailed individual spectral modeling needed to calculate accurate stellar masses of our galaxies is beyond the scope of this work, we obtained rough estimates for each of the galaxy sub-groups in our study. We base these on the integrated i-band light given in the SDSS photometric database (https://www.sciserver.org/, using DR7 to match our spectroscopic data). Since the average redshifts of all of our AGN sub-groups are the same, it is not surprising that their average total i-band luminosities, assumed to be almost entirely starlight, are also very similar. In fact, the average absolute magnitude of all categories of Seyfert 1’s, Sy2’s and LINERs are within 0.1 mag of -21.6 (Table 4). Assuming the mass/light ratio of typical early type galaxies, of 2.5 in solar units, this corresponds to . We obtain similar stellar masses for our “Lower” and “Middle” SF galaxies, which have mean absolute i magnitude of -21.1 and -20.8 respectively. (The average absolute magnitude of our SF “Lower” galaxies, which we argue do not match the higher metallicities of Seyfert host galaxies, is -19.8).
IV Predicting [NII] Flux from [SII] to Deblend Halpha
@@
V Reddening Estimates from Narrow Emission Line Ratios
@@ V.1 Milky Way Reddening
V.2 Average Reddenings of AGN Categories
@@
V.3 Comparison of Balmer Decrements with [SII]/[OII] Ratios
@@
VI Electron Densities in Different Galaxy Classes
@@
VII Are Weak Seyfert 1’s Explained by Low Ionization Parameter, Rather than Star Formation Contamination?
@@
VIII Absorption Line Measurements of Stellar Populations in Seyfert Galaxies
@@
IX Small Contamination from Seyfert 1 Nonstellar Light
@@
The dashed line in each of the Figures 12a–d is a simple straight-line fit for pure starlight being emitted from pure stellar populations in a galaxy, from the top where SF-Lower galaxies are located, down to the LINERs in the lower right of these figures. This sequence is based on the fact that the continuum in star-forming galaxies is dominated by light from hot, luminous young stars, whereas LINER galaxies contain an older stellar population with no star formation in the last Gyr.
@@
We searched further for evidence of nonstellar AGN light by examining two additional measures of the continuum which are sensitive to violet light, where the AGN contribution is strongest. We show in Figure 12d the integrated color of the entire galaxy, obtained directly from the SDSS SkyServer photometry catalog, averaged over each galaxy class. In Figure 12, both and D4000 are compared against our stellar population age indicator, H absorption. The same stellar population age sequence is evident. Again, the color and especially D4000 become smaller in the less AGN-dominated spectra. In other words, the intermediate Sy1’s, the Sy2L’s, and especially the weak Sy1’s show this indication of a larger proportion of blue continuum light from younger stars. Their average stellar populations are significantly younger than those of Passive Galaxies, LINERs, and Sy2U’s.
However, there is an additional effect evident in the color index: the strong Sy1’s (Sy1.2 and especially Sy1.0) shift dramatically to the blue (left), even though their weak H absorption indicates few young stars. Thus the strong Seyfert 1’s have substantial extra flux in the u bandpass, which is not produced by hot stars. Instead, we attribute this extra u light to strong BLR emission contamination. Indeed, M. A. Malkan & W. L. W. Sargent (1982) first measured the BLR contributions produced by the blended broad-lines of the higher Balmer series above H, which then merge into a pseudo-continuum rising smoothly up to 3646Å, the onset of the Balmer Continuum emission jump. They found that the Balmer jump can produce as much flux at 3646Å as the underlying continuum. Based on their finding that the Balmer continuum flux is relatively flat throughout most of the sensitive part of the u filter bandpass (3350–3750Å), the typical Sy1 studied by M. A. Malkan & W. L. W. Sargent (1982) would have its integrated u flux enhanced by a factor of roughly 50%. The horizontal vector shows the effect of this BLR contamination on shifting the color of strong Sy1’s up to 0.44 mag to the left. In the intermediate and weak Seyfert 1’s, with substantially less broad H emission, this BLR contamination is too small to be measurable.
@@
X Selection Effects in Seyfert Galaxy Surveys
@@
X.1 “Strong Unification” of Seyfert 1 and 2 Galaxies
@@
X.2 Are All Broad Line Regions the Same?